/ means square root.
[] means absoulte value..
1.. simpliy -/-169
2..find [-1+i]




Answer by stanbon(75887)

/ means square root.
[] means absolute value..
1.. simpliy -/-169
-sqrt(-169) = -13i
------------------------------
2..find [-1+i]
|-1+i| = sqrt((-1)^2 + 1^2)
= sqrt(2)
=============
Cheers,
Stan H.